### v2

# Text-to-Vision Generation with Scene Graph Programming

Abstract
Generative models like DALL-E and Sora have gained attention by producing implausible images, such as “astronauts riding a horse in space.” Despite the proliferation of text-to-vision models that have inundated the internet with synthetic visuals, from images to 3D assets, current benchmarks predominantly evaluate these models on real-world scenes paired with captions. We introduce Generate Any Scene, a framework that systematically enumerates scene graphs representing a vast array of visual scenes, spanning realistic to imaginative compositions. Generate Any Scene leverages ‘scene graph programming,’ a method for dynamically constructing scene graphs of varying complexity from a structured taxonomy of visual elements. This taxonomy includes numerous objects, attributes, and relations, enabling the synthesis of an almost infinite variety of scene graphs. Using these structured representations, Generate Any Scene translates each scene graph into a caption, enabling scalable evaluation of text-to-vision models through standard metrics. We conduct extensive evaluations across multiple text-to-image, text-to-video, and text-to-3D models, presenting key findings on model performance. We find that DiT-backbone text-to-image models align more closely with input captions than UNet-backbone models. Text-to-video models struggle with balancing dynamics and consistency, while both text-to-video and text-to-3D models show notable gaps in human preference alignment. Additionally, we demonstrate the effectiveness of Generate Any Scene by conducting three practical applications leveraging captions generated by Generate Any Scene: (1) a self-improving framework where models iteratively enhance their performance using generated data, (2) a distillation process to transfer specific strengths from proprietary models to open-source counterparts, and (3) improvements in content moderation by identifying and generating challenging synthetic data.
1 Introduction
Figure 1: Overview of applications with Generate Any Scene captions: Application 1: (Self-improving): Iteratively enhances a model by generating images with Generate Any Scene captions, selecting the best, and fine-tuning, yielding a performance boost. Application 2: (Distilling limitations): Distills strengths from proprietary models, such as better compositionality and hard concept understanding, into open-source models. Application 3: (Generated content detector): Robustify AI-generated content detection by training on diverse synthetic data generated by Generate Any Scene’s captions.
Artist Marc Chagall said “Great art picks up where nature ends.” The charm of visual content generation lies in the realm of imagination. Since their launch, Dall-E [55, 5] and Sora [7] have promoted their products with implausible generated images of “astronauts riding a horse in space” and “cats playing chess”. With the proliferation of text-to-vision generation models, the internet is now flooded with generated visual content—images, videos, and 3D assets—most generated from user-provided captions [55, 5, 7]. While there are numerous benchmarks designed for evaluating these text-to-vision models, they are typically collections of real-world visual content paired with captions [9, 34, 70]. To quote Marc Chagall again, “If I create from heart, nearly everything works; if from the head, almost nothing.” There is a need for evaluation benchmarks that go beyond real-world scenes and evaluate how well generative models can represent the entire space of imaginary scenes.
Such a comprehensive evaluation requires that we first define the space of the visual content. A long list of prior work [32, 27, 47, 28, 26] has argued that scene graphs [32] are a cognitively grounded [6] representation of the visual space. A scene graph represents objects in a scene as individual nodes in a graph. Each object is modified by attributes, which describe its properties. For example, attributes can describe the material, color, size, and location of the object in the scene. Finally, relationships are edges that connect the nodes. They define the spatial, functional, social, and interactions between objects [42]. For example, in a living room scene, a “table” node might have attributes like “wooden” or “rectangular” and be connected to a “lamp” node through a relation: “on top of.” This systematic scene graph structure provides simple yet effective ways to define and model the scene. Make it an ideal structure for Generate Any Scene to systematically define the diverse space of the visual scenes.
We introduce Generate Any Scene, a system capable of efficiently enumerating the space of scene graphs representing a wide range of visual scenes, from realistic to highly imaginative. Generate Any Scene is powered by what we call scene graph programming, a programmatic approach for composing scene graphs of any complexity using a rich taxonomy of visual elements, and for translating each scene graph into a caption. With a space of synthetically diverse captions, we use Generate Any Scene to prompt Text-to-Vision generation models and evaluate their generations. Like any other representation, scene graphs are also limited: they don’t represent tertiary relationships (e.g. “three people playing frisbee”). Nonetheless, they account for a large space of possibilities. To systematically define and scalably explore the space of user captions, we adopt the scene graph representation [32] to comprehensively evaluate and improve text-to-vision models.
We construct a rich taxonomy of visual concepts consisting of $28,787$ objects, $1,494$ attributes, $10,492$ relations, $2,193$ image/video/3D scene attributes from various sources. Based on these assets, Generate Any Scene can programmatically synthesize an almost infinite number of scene graphs of varying complexity [81]. Besides, Generate Any Scene allows configurable scene graph generation. For example, evaluators can specify the complexity level of the scene graph to be generated or provide a seed scene graph to be expanded. Given an initial scene graph, Generate Any Scene programmatically translates it into a caption, which, when combined with existing text-to-vision metrics, e.g., Clip Score [54] and VQA Score [39], can be used to evaluate any text-to-vision model [62]. By automating these steps, our system ensures both scalability and adaptability, providing researchers and developers with diverse, richly detailed scene graphs and corresponding captions tailored to their specific needs.
With Generate Any Scene’s programmatic generation capability, we release a dataset featuring $10$ million diverse and compositional captions, each paired with a corresponding scene graph. This extensive dataset spans a wide range of visual scenarios, from realistic to highly imaginative compositions, providing an invaluable resource for researchers and practitioners in the Text-to-Vision generation field. We also conduct extensive evaluations of $12$ text-to-image, $9$ text-to-video and $5$ text-to-3D models across a broad spectrum of visual scenes. We have several crucial findings: (1) DiT-backbone models show superior faithfulness and comprehensiveness to input captions than UNet-backbone models, with human-alignment data training helping to bridge some of these gaps. (2) Text-to-Video generation face challenges in balancing dynamics and consistency. (3) All Text-to-Video and Text-to-3D models we evaluate show negative ImageReward Score scores, highlighting a substantial gap in human preference alignment.
Further, we demonstrate the effectiveness of Generate Any Scene by conducting three practical applications leveraging captions generated by Generate Any Scene (Figure 1):
Application 1: Self-improving. We show that our diverse captions can facilitate a framework to iteratively improve Text-to-Vision generation models using their own generations. Given a model, we generate multiple images, identify the highest-scoring one, and use it as new fine-tuning data to improve the model itself. We fine-tune Stable Diffusion v1-5 and achieve an average of 5% performance boost compared with original models, and this method is even better than fine-tuning with the same amount of real images and captions from the Conceptual Captions CC3M [9] over different benchmarks.
Application 2: Distilling limitations. Using our evaluations, we identify limitations in open-sourced models that their proprietary counterparts excel at. Next, we distill these specific capabilities from proprietary models. For example, DaLL-E 3 excels particularly in generating composite images with multiple parts. We distill this capability into Stable Diffusion v1-5, effectively bridging the gap between DaLL-E 3 and Stable Diffusion v1-5.
Application 3: Generated content detector. Content moderation is a vital application, especially as Text-to-Vision generation models improve. We identify which kinds of data content moderation models are bad at detecting, generate more of such content, and retrain the detectors. We train a ViT-T with our generated data and boost its detection capabilities across benchmarks.
2 Related work
Text-to-Vision generation models.
Recent Text-to-Image generation advances are driven by diffusion models that enhance visual fidelity and semantic alignment. Some open-source models [56, 52, 35, 51, 1] use UNet backbones to refine images iteratively. In parallel, Diffusion Transformers (DiTs) architectures[16, 11, 12, 33] have emerged as a better alternative in capturing long-range dependencies and improving coherence. Proprietary models like DALL-E 3 [5] and Imagen 3 [2] still set the state-of-the-art. Based on Text-to-Image generation method, Text-to-Video generation models typically utilize time-aware architectures to ensure temporal coherence across frames [19, 65, 30, 67, 74, 10, 80, 84]. In Text-to-3D generation, recent proposed models [69, 53, 66, 44, 38] integrate the diffusion models with Neural Radiance Fields (NeRF) rendering to generate diverse 3D object. In this work, we systematically evaluate and deeply analyze these Text-to-Vision generation models.
Synthetic prompts for Text-to-Vision generation.
Prompts for Text-to-Vision generation models vary greatly in diversity, complexity, and compositionality. This variation makes it challenging and costly to collect large-scale and diverse prompts written by humans. Consequently, synthetic prompts have been widely used for both training [37, 62, 36, 83, 43, 63, 48, 71, 76, 82] and evaluation purposes [23]. For example, training methods like LLM-Grounded Diffusion [37] leverage LLM-generated synthetic text prompts to enhance the model’s understanding and alignment with human instruction. For evaluation, benchmarks such as T2I-CompBench [23] and T2V-CompBench [63] utilize benchmarks generated by LLMs. However, while LLMs can generate large-scale natural prompts, controlling their content is difficult, and they might exhibit systematic bias. In this work, we propose a programmatic scene graph-based prompt generation system that can generate infinitely diverse scene graph-based prompts for evaluating and improving Text-to-Vision generation models.
Finetuning techniques for Text-to-Vision generation.
To accommodate the diverse applications and personalization needs in text-to-vision models, numerous fine-tuning techniques have been developed. For example, LoRA [21] reduces the computational resources required for fine-tuning by approximating weight updates with low-rank matrices, enabling efficient adaptation to new tasks. Textual Inversion [46, 17] introduces new word embeddings to represent novel concepts, allowing models to generate images of user-specified content without extensively altering the original parameters. DreamBooth [57] fine-tunes models on a small set of personalized images to capture specific subjects or styles, facilitating customized content generation. DreamSync [62], provides a novel way to enable models to learn from the high-quality output generated by itself. In this work, we use several fine-tuning techniques with our programmatic scene graph-based prompts to improve Text-to-Vision generation models.
3 Generate Any Scene
Figure 2: The generation pipeline of Generate Any Scene: Step 1: The system enumerates scene graph structures that contain objects, attributes, and relations based on complexity, and queries the corresponding scene graph structure that satisfies the needs. Step 2: It populates these structures with metadata, assigning specific content to each node. Scene graphs are completed in this step. Step 3: In addition to the scene graph, scene attributes—such as art style and camera settings—are sampled to provide contextual depth beyond the scene graph. Step 4: The Generate Any Scene system combines the scene graph and scene attributes, such as art style and camera settings, and then translates them into a coherent caption by organizing the elements into structured text.
We present our implementation of Generate Any Scene system. (Figure 2) It programmatically synthesizes diverse scene graphs in terms of both structure and content and translates them into corresponding captions.
Scene graph. A scene graph is a structured representation of a visual scene, where objects are represented as nodes, their attributes (such as color and shape) are properties of those nodes, and their relationships (such as spatial or semantic connections) are represented as edges. In recent years, scene graphs have played a crucial role in visual understanding tasks, such as those found in Visual Genome [32] and GQA [25] for visual question answering (VQA). Their utility has expanded to various Text-to-Vision generation tasks. For example, the DSG score [13] leverages MLMs to evaluate how well captions align with generated scenes by analyzing scene graphs.
Metadata Type Number Source
Objects 28,787 WordNet [45]
Attributes 1,494 Wikipedia [72], etc.
Relations 10,492 Robin [49]
Scene Attributes 2,193 Places365 [41], etc.
Table 1: Summary of the quantities and source of visual elements.
Taxonomy of visual elements. To construct a scene graph, we use three main metadata types: objects, attributes, and relations. We also have scene attributes that capture the board aspect of the caption, such as art style, to create a complete visual caption. The numbers and the source of our metadata are illustrated in Table 1. Additionally, we build a taxonomy that categorizes metadata into distinct levels and types, enabling fine-grained analysis. This structure allows for detailed assessments, such as evaluating model performance on “flower” as a general concept and on specific subcategories like “daisy.” More details in Appendix C.
3.1 Scene graph programming
Step 1: Scene graph structure enumeration and query. Our system first generates and stores a variety of scene graph structures based on a specified level of complexity, defined by the total number of objects, relationships, and attributes in each graph. The process begins by determining the number of object nodes, and then by systematically enumerating different combinations of relationships among these objects and their associated attributes. Once all graph structures meeting the complexity constraint are enumerated, they are stored in a database for later use. This enumeration process is executed only once for each level of complexity, allowing us to efficiently query the database for suitable templates when needed.
Step 2: Populate the scene graph structure with metadata. Given a scene graph structure, the next step involves populating the graph with metadata. For each object node, attribute node, and relation edge, we sample the corresponding content from our metadata. This process is highly customizable: users can define the topics and types of metadata to be included (e.g., selecting only common metadata or specifying particular relationships between particular objects, among other options). By determining the scope of metadata sampling, we can precisely control the final content of the captions and easily extend the diversity and richness in the scene graphs by incorporating new datasets.
Step 3: Sampling scene attributes. In addition to scene graphs that capture the visual content of the image, we also include scene attributes that describe aspects such as the art style, viewpoint, time span (for video), and 3D attributes (for 3D content). These scene attributes are sampled directly from our metadata, creating a list that provides contextual details to enrich the description of the visual content.
Step 4: Translate scene graph to caption. We introduce an algorithm that converts scene graphs and a list of scene attributes into captions. The algorithm processes the scene graph in topological order, transforming each object, its attributes, and relational edges into descriptive text. To maintain coherence, it tracks each concept’s occurrence, distinguishing objects with identical names using terms like “the first” or “the second.” Objects that have been previously referenced without new relations are skipped to avoid misreferencing. This approach enhances caption clarity by preventing repetition and maintaining a logical reference.
4 Evaluating Text-to-Vision generation models
4.1 Experiment Settings
Details of experiment settings are in Appendix D.
Models. We conduct experiments on 12 Text-to-image models [1, 52, 56, 35, 51, 16, 11, 12, 33, 5], 9 Text-to-Video models [67, 61, 30, 19, 65, 74, 84, 80, 10], and 5 Text-to-3D models [53, 38, 66, 69, 44]. Text-to-image models are evaluated at a resolution of 1024 × 1024 pixels. We standardize the frame length to 16 across all Text-to-Video models for fair comparisons. For Text-to-3D, we generate videos by rendering from 120 viewpoints.
Metrics. Across all Text-to-Vision generation tasks, we use Clip Score [8] (semantic similarity), VQA Score [39] (faithfulness), TIFA Score [22, 13] (faithfulness), Pick Score [31] (human preference), and ImageReward Score [77] (human preference) as general metrics, and for Text-to-Video generation, VBench [24] for fine-grained video analysis like consistency and dynamics.
Synthetic captions. We evaluate our Text-to-Image generation and Text-to-Video generation models on 10K randomly generated captions, with scene graph complexity ranging from 3 to 12 and scene attributes from 0 to 5, using unrestricted metadata. For Text-to-3D generation models, due to their limitations in handling complex captions and time-intensive generation, we restrict scene graph complexity to 1-3, scene attributes to 0-2, and evaluate on 1K captions.
4.2 Overall results
We evaluate Text-to-Image generation, Text-to-Video generation, and Text-to-3D generation models on Generate Any Scene. Here, we only list key findings; more details and raw results can be found in Appendix D.
Text-to-Image generation results. (Figure 3)
Figure 3: Comparative evaluation of Text-to-Image generation models across different backbones (DiT and UNet) using multiple metrics: TIFA Score, Pick Score, VQA Score, and ImageReward Score.
  1. 1.
    DiT-backbone models outperform UNet-backbone models on VQA Score and TIFA Score, indicating greater faithfulness and comprehensiveness to input captions.
  2. 2.
    Despite using a UNet architecture, Playground v2.5 achieves higher Pick Score and ImageReward Score scores than other open-source models. We attribute this to Playground v2.5 ’s alignment with human preferences achieved during training.
  3. 3.
    The closed-source model DaLL-E 3 maintains a significant lead in VQA Score, TIFA Score, and ImageReward Score, demonstrating strong faithfulness and alignment with prompts across generated content.
Text-to-Video generation results. (Table 2,3)
Model clip score pick score image reward score VQA score TiFA score
VideoCraft2 [10] 0.2398 0.1976 -0.4202 0.5018 0.2466
AnimateLCM [65] 0.2450 0.1987 -0.5754 0.4816 0.2176
AnimateDiff [19] 0.2610 0.1959 -0.7301 0.5255 0.2208
Open-Sora 1.2 [84] 0.2259 0.1928 -0.6277 0.5519 0.2414
FreeInit [74] 0.2579 0.1950 -0.9335 0.5123 0.2047
ModelScope [67] 0.2041 0.1886 -1.9172 0.3840 0.1219
Text2Video-Zero [30] 0.2539 0.1933 -1.2050 0.4753 0.1952
CogVideoX-2B [80] 0.2038 0.1901 -1.2301 0.4585 0.1997
ZeroScope [61] 0.2289 0.1933 -1.1599 0.4892 0.2388
Table 2: Overall performance of Text-to-Video generation models over 10K Generate Any Scene captions.Red Cell is the highest score. Yellow Cell is the second highest score.
Model subject consistency background consistency motion smoothness dynamic degree
Open-Sora 1.2 0.9964 0.9907 0.9973 0.0044
Text2Video-Zero 0.8471 0.9030 0.8301 0.9999
VideoCraft2 0.9768 0.9688 0.9833 0.3556
AnimateDiff 0.9823 0.9733 0.9859 0.1406
FreeInit 0.9581 0.9571 0.9752 0.4440
ModelScope 0.9795 0.9831 0.9803 0.1281
AnimateLCM 0.9883 0.9802 0.9887 0.0612
CogVideoX-2B 0.9583 0.9602 0.9823 0.4980
ZeroScope 0.9814 0.9811 0.9919 0.1670
Table 3: Overall performance of Text-to-Video generation models over 10K Generate Any Scene captions with VBench metrics. Red Cell is the highest score. Blue Cell is the lowest score.
  1. 1.
    Text-to-video models face challenges in balancing dynamics and consistency (Table 3). This is especially evident in Open-Sora 1.2, which achieves high consistency but minimal dynamics, and Text2Video-Zero, which excels in dynamics but suffers from frame inconsistency.
  2. 2.
    All models exhibit negative ImageReward Score (Table 2), suggesting a lack of human-preferred visual appeal in the generated content, even in cases where certain models demonstrate strong semantic alignment.
  3. 3.
    VideoCrafter2 strikes a balance across key metrics, leading in human-preference alignment, faithfulness, consistency, and dynamic.
Text-to-3D generation results. (Table 4)
Model clip score pick score vqa score tifa score image reward score
Latent-NeRF [44] 0.2115 0.1910 0.4767 0.2216 -1.5311
DreamFusion-sd [53] 0.1961 0.1906 0.4421 0.1657 -1.5582
Magic3D-sd [38] 0.1947 0.1903 0.4193 0.1537 -1.6327
SJC [66] 0.2191 0.1915 0.5015 0.2563 -1.4370
DreamFusion-IF [53] 0.1828 0.1857 0.3872 0.1416 -1.9353
Magic3D-IF [38] 0.1919 0.1866 0.4039 0.1537 -1.8465
ProlificDreamer [69] 0.2125 0.1940 0.5411 0.2704 -1.2774
Table 4: Overall performance of Text-to-3D generation models over 10K Generate Any Scene captions.
  1. 1.
    ProlificDreamer outperforms other models, particularly in ImageReward Score, VQA Score and TIFA Score.
  2. 2.
    All models receive negative ImageReward Score scores, highlighting a significant gap between human preference and current Text-to-3D generation generation capabilities.
5 Application 1: Self-Improving Models
Figure 4: Results for Application 1: Self-Improving Models. Average VQA score of Stable Diffusion v1-5 fine-tuned on different data across 1K Generate Any Scene image/video evaluation set and GenAI-Bench image/video benchmark [34].
In this section, we explore how Generate Any Scene facilitates a self-improvement framework for model generation capabilities. By programmatically generating scalable compositional captions from scene graphs, Generate Any Scene expands the textual and visual space, allowing for a diversity of synthetic images that extend beyond real-world scenes. Our goal is to utilize these richly varied synthetic images to further boost model performance.
Iterative self-improving framework. Inspired by DreamSync [62], we designed an iterative self-improving framework using Generate Any Scene with Stable Diffusion v1-5 as the baseline model. With VQA Score, which shows strong correlation with human evaluations on compositional images [39], we guide the model’s improvement throughout the process.
Specifically, Generate Any Scene generates 3 $\times$ 10K captions across three epochs. For each caption, Stable Diffusion v1-5 generates 8 images, and the image with the highest VQA Score is selected. From each set of 10K optimal images, we then select the top 25% (2.5k image-caption pairs) as the training data for each epoch. In subsequent epochs, we use the fine-tuned model from the prior iteration to generate new images. We employ LoRA [21] for parameter-efficient fine-tuning. Additional details are available in Appendix E.
To evaluate the effectiveness of self-improvement using synthetic data generated by Generate Any Scene, we conduct comparative experiments with the CC3M dataset, which comprises high-quality and diverse real-world image-caption pairs [60]. We randomly sample 3 $\times$ 10K captions from CC3M, applying the same top-score selection strategy for iterative fine-tuning of Stable Diffusion v1-5. Additionally, we include a baseline using random-sample fine-tuning strategy to validate the advantage of our highest-scoring selection-based strategy.
Results. We evaluate our self-improving pipeline on Text-to-Vision generation benchmarks, including GenAI Bench [34]. For the Text-to-Video generation task, we use Text2Video-Zero as the baseline model, substituting its backbone with the original Stable Diffusion v1-5 and our fine-tuned Stable Diffusion v1-5 models.
Our results show that fine-tuning with Generate Any Scene-generated synthetic data consistently outperforms CC3M-based fine-tuning across Text-to-Vision generation tasks (Figure 4), achieving the highest gains with our highest-scoring selection strategy. This highlights Generate Any Scene’s scalability and compositional diversity, enabling models to effectively capture complex scene structures. Additional results are in Appendix F.
Takeaway for application 1 Iterative self-improving Text-to-Vision generation models with compositional and diverse synthetic captions can surpass fine-tuning with real-world image-caption data. Potential reason: The compositional, synthetic captions generated by Generate Any Scene exhibit greater diversity than real-world data.
6 Application 2: Distilling limitations
Although self-improving with Generate Any Scene-generated data shows clear advantages over high-quality real-world datasets, its efficiency remains inherently constrained by the limitations of the model’s own generation ability. To address this, we leverage the taxonomy and programmatic generation capabilities within Generate Any Scene to identify specific strengths of proprietary models (DaLL-E 3), and to distill these capabilities into open-source models. More details are in Appendix F.
6.1 Fine-Grained Analysis of DaLL-E 3’s Exceptional Performance
Figure 5: Examples for Application 2: Distilling limitations. Examples of images generated by DaLL-E 3, the original Stable Diffusion v1-5, and the fine-tuned versions. The left four captions demonstrate fine-tuning with multi-object captions generated by Generate Any Scene for better compositionality, while the right two columns focus on understanding hard concepts.
As shown in Figure 3, DaLL-E 3 achieves TIFA Score 1.5 to 2 times higher than those of other models. When we compare TIFA Score across varying numbers of elements (objects, relations, and attributes per caption) in Figure 6(b), DaLL-E 3 counterintuitively maintains consistent performance regardless of element count. The performance of other models declines as the element count increases, which aligns with expected compositional challenges. We suspect these differences are primarily due to DaLL-E 3’s advanced capabilities in compositionality and understanding hard concepts, which ensures high faithfulness across diverse combinations of element types and counts.
6.2 Distilling compositionality from DaLL-E 3
Observations. We find that DaLL-E 3 tends to produce straightforward combinations of multiple objects (Figure 5). In contrast, open-source models like Stable Diffusion v1-5 often omit some objects from the captions, even though they are capable of generating each object individually.
This difference suggests that DaLL-E 3 may be trained on datasets emphasizing multi-object presence without rigorous attention to image layout or object interaction. Such training likely underpins DaLL-E 3’s stronger performance on metrics like TIFA Score and VQA Score, prioritizing object inclusion over detailed compositional arrangement.
Finetuning. To encourage Stable Diffusion v1-5 to learn compositional abilities similar to those of DaLL-E 3., we select a set of 778 images generated by DaLL-E 3, each containing multiple objects, and utilize this dataset to fine-tune Stable Diffusion v1-5. For the baseline, we randomly sampled an equivalent set of DaLL-E 3-generated images paired with generated captions from Generate Any Scene.
Results. To evaluate compositional improvements, we generate 1K multi-object captions. Figure 6(b) shows a 10% TIFA Score increase after fine-tuning, random fine-tuning by an average of 3%. These results indicate enhanced compositional abilities in handling complex generation tasks.
We analyze images generated by Stable Diffusion v1-5 before and after fine-tuning on high-complexity image-caption pairs (Figure 5). It is surprising to see that, with only 1K LoRA fine-tuning steps, Stable Diffusion v1-5 effectively learn DaLL-E 3 ’s capability to arrange and compose multiple objects within a single image,. This fine-tuning strategy notably enhances alignment between generated images and their given captions.
On a broader set of 10K Generate Any Scene-generated captions, the fine-tuned model consistently outperformed the randomly fine-tuned model (Figure 6(a)), confirming the generalizability and superiority of targeted fine-tuning for improving model performance.
(a) Distilling compositionality from DaLL-E 3: Model results on TIFA vs. total element numbers in captions in 10K general Generate Any Scene captions.
(b) Distilling compositionality from DaLL-E 3: Model results on TIFA vs. total element numbers in captions in 1K multi-object Generate Any Scene captions.
(c) Learning hard concepts from DALL-E 3: Models’ average TIFA Score performance over captions and hard concepts in 1K hard concepts Generate Any Scene captions.
Figure 6: Results for Application 2: Distilling limitations. The left two figures show the results for Distilling compositionality from DALL-E 3, while the rightmost figure shows the results for Learning hard concepts from DALL-E 3.
6.3 Learning hard concepts from DaLL-E 3
Observation. Figure 5 shows that is capable not only of handling multi-object generation but also of understanding and generating rare and hard concepts, such as a specific species of flower. We attribute this to its training with proprietary real-world data.
Finetuning. Using the taxonomy of Generate Any Scene, we compute model performance on each conceptby averaging scores across captions containing that concept.Accumulating results through the taxonomy, we identify the 100 concepts where Stable Diffusion v1-5 shows the largest performance gap relative to DaLL-E 3. For fine-tuning, we generate 778 captions incorporating these concepts with others, using DaLL-E 3 to produce corresponding images. As a baseline, we randomly select 778 Generate Any Scene-generated captions for fine-tuning and compare these with the original Stable Diffusion v1-5 model.
(a) In-domain testing (Same Model - SD v1.4): Detection results on images generated by SD v1.4 using the GenImage dataset.
(b) In domain testing (cross-model):Average detection results on images generated by multiple models using our captions.
(c) Out of domain: Average detection results on images generated by multiple models using captions from the GenImage dataset.
Figure 7: Results for Application 3: Generated content detector. Comparison of detection performance across different data scales using $D^{3}$ alone versus the combined $D^{3}$ + Generate Any Scene training set in cross-model and cross-dataset scenarios.
Results. The results in Figure 6(c) show that our targeted fine-tuning led to improved model performance, reflected in higher average scores across captions and increased scores for each challenging concept.
Takeaway for application 2 Targeted fine-tuning can distill proprietary model strengths, effectively bridging gaps in compositionality and concept handling for open-source models. Potential Reason: Generate Any Scene facilitates fine-grained analysis to identify specific performance gaps, enabling targeted data selection to distill limitations.
7 Application 3: Generated content detector
Advances in Text-to-Vision generation underscore the need for effective content moderation [50]. Major challenges include the lack of high-quality and diverse datasets and the difficulty of generalizing detection across models Text-to-Vision generation [68, 29]. Generate Any Scene addresses these issues by enabling scalable, programmatic generation of compositional captions, increasing the diversity and volume of synthetic data. This approach enhances existing datasets by compensating for their limited scope-from realistic to imaginative-and variability.
To demonstrate Generate Any Scene’s effectiveness in training generated content detectors, we used the $D^{3}$ dataset [3] as a baseline. We sampled 5k captioned real and SDv1.4-generated image pairs from $D^{3}$ and generated 5k additional images with Generate Any Scene captions. We trained a ViT–T [73] model with a single-layer linear classifier, varying dataset sizes with $N$ real and $N$ synthetic images. For synthetic data, we compared $N$ samples solely from $D^{3}$ with a mixed set of $N/2$ from Generate Any Scene and $N/2$ from $D^{3}$, keeping the same training size.
We evaluate the detector’s generalization on the GenImage [85] validation set and images generated using Generate Any Scene captions. Figure 7 demonstrates that combining Generate Any Scene-generated images with real-world captioned images consistently enhances detection performance, particularly across cross-model scenarios and diverse visual scenes. More details are in Appendix G.
Takeaway for application 3 Compositional synthetic captions robustify generated content detectors. Potential reason: Generate Any Scene can generate more diverse captions to complement real-world image-caption training data by enriching compositional variety and imaginative scope.
8 Limitation
Programmatically generated prompts can be unrealistic and biased.
Programmatically generated prompts can be unrealistic and biased. Although our system is capable of producing a wide range of rare compositional scenes and corresponding prompts, some of these outputs may violate rules or conventions, going beyond what is even considered imaginable or plausible. We also implement a pipeline to filter the commonsense of the generated prompts using the Vera score (a large language model-based commonsense metric) and Perplexity, but we make this pipeline optional.
Linguistic diversity of programmatic prompts is limited.
While Generate Any Scene excels at generating diverse and compositional scene graphs and prompts, its ability to produce varied language expressions is somewhat constrained. The programmatic approach to generating content ensures diversity in terms of the elements of the scene, but it is limited when it comes to linguistic diversity and the richness of expression. To address this, we introduce a pipeline that leverages large language models (LLMs) to paraphrase prompts, enhancing linguistic variety. However, this addition introduces new challenges. LLMs are prone to biases and hallucinations, which can affect the quality and reliability of the output. Furthermore, the use of LLMs risks distorting the integrity of the original scene graph structure, compromising the coherence and accuracy of the generated content. So we make this LLM paraphrase pipeline optional for our paper.
9 Conclusion
We present Generate Any Scene, a system leveraging scene graph programming to generate diverse and compositional synthetic captions for Text-to-Vision generation tasks. It extends beyond existing real-world caption datasets to include imaginary scenes and even implausible scenarios. To demonstrate the effectiveness of Generate Any Scene, we explore three applications: (1) self-improvement by iteratively optimizing models, (2) distillation of proprietary model strengths into open-source models, and (3) robust content moderation with diverse synthetic data. Generate Any Scene highlights the importance of synthetic data in evaluating and improving Text-to-Vision generation, and addresses the need to systematically define and scalably produce the space of visual scenes.
Acknowledgement
This project was partially supported by an OpenAI “Superalignment Fast” grant.
Supplementary Material
Appendix A More Analysis with Generate Any Scene
With Generate Any Scene, we can generate infinitely diverse and highly controllable prompts. Using Generate Any Scene, we conduct several analyses to provide insights into the performance of today’s Text-to-Vision generation models.
A.1 Performance analysis across caption properties
In this section, we delve into how model performance varies with respect to distinct properties of Generate Any Scene captions. While Generate Any Scene is capable of generating an extensive diversity of captions, these outputs inherently differ in key characteristics that influence model evaluation. Specifically, we examine three properties of the caption: Commonsense, Perplexity, and Scene Graph Complexity (captured as the number of elements in the captions). These properties are critical in understanding how different models perform across a spectrum of linguistic and semantic challenges presented by captions with varying levels of coherence, plausibility, and compositional richness.
Perplexity. (Figure 8)
Figure 8: Average performance of models across different percentiles of perplexity of captions, evaluated on various metrics. From left to right, the perplexity decreases, indicating captions that are progressively more reasonable and easier for the LLM to generate.
Perplexity is a metric used to measure a language model’s unpredictability or uncertainty in generating a text sequence. A higher perplexity value indicates that the sentences are less coherent or less likely to be generated by the model.
As shown in Figure 8, From left to right, when perplexity increases, indicating that the sentences become less reasonable and less typical of those generated by a language model, we observe no clear or consistent trends across all models and metrics. This suggests that the relationship between perplexity and model performance varies depending on the specific model and evaluation metric.
Commonsense. (Figure 9)
Figure 9: Average performance of models across different percentiles of Vera Score for captions, evaluated on various metrics. From left to right, the Vera Score decreases, indicating captions that exhibit less commonsense reasoning and are more likely to describe implausible scenes.
Figure 10: Average performance of models across different numbers of elements (objects + attributes + relations) in the scene graph (complexity of the scene graph) of the captions, evaluated on various metrics. From left to right, as the number of elements (complexity) increases, the scene graphs become more complicated and compositional.
Commonsense is an inherent property of text. We utilize the Vera Score [40], a metric generated by a fine-tuned LLM to evaluate the text’s commonsense level.
As shown in Figure 9, from left to right, as the Vera Score increases—indicating that the captions exhibit greater commonsense reasoning—we observe a general improvement in performance across all metrics and models, except for Clip Score. This trend underscores the correlation between commonsense-rich captions and enhanced model performance.
Element Numbers (Complexity of Scene Graph). (Figure 10)
Finally, we evaluate model performance across total element numbers in the captions, which represent the complexity of scene graphs (objects + attributes + relations).
From left to right, the complexity of scene graphs becomes higher, reflecting more compositional and intricate captions. Across most metrics and models, we observe a noticeable performance decline as the scene graphs become more complex. However, an interesting exception is observed in the performance of DaLL-E 3. Unlike other models, DaLL-E 3 performs exceptionally well on VQA Score and TIFA Score, particularly on VQA Score, where it even shows a slight improvement as caption complexity increases. This suggests that DaLL-E 3 may have a unique capacity to handle complex and compositional captions effectively.
A.2 Analysis on different metrics
Compared with most LLM and VLM benchmarks that use multiple-choice questions and accuracy as metrics. There is no universal metric in evaluating Text-to-Vision generation models. Researchers commonly used model-based metrics like Clip Score, VQA Score, etc. Each of these metrics is created and fine-tuned for different purposes with bias. Therefore, we also analysis on different metrics.
Clip Score isn’t a universal metric.
Clip Score is one of the most widely used metrics in Text-to-Vision generation for evaluating the alignment between visual content and text. However, our analysis reveals that Clip Score is not a perfect metric and displays some unusual trends. For instance, as shown in Figures 8, 9, and 10, we compute the perplexity across 10k prompts used in our study, where higher perplexity indicates more unpredictable or disorganized text. Interestingly, unlike other metrics, Clip Score decreases as perplexity lowers, suggesting that Clip Score tends to favor more disorganized text. This behavior is counterintuitive and highlights the potential limitations of using Clip Score as a robust alignment metric.
Limitations of human preference-based metrics.
We use two metrics fine-tuned using human preference data: Pick Score and ImageReward Score. However, we found that these metrics exhibit a strong bias toward the data on which they were fine-tuned. For instance, as shown in Table 5, Pick Score assigns similar scores across all models, failing to provide significant differentiation or meaningful insights into model performance. In contrast, ImageReward Score demonstrates clearer preferences, favoring models such as DaLL-E 3 and Playground v2.5, which incorporated human-alignment techniques during their training. However, this metric shows a significant drawback: it assigns disproportionately large negative scores to models like Stable Diffusion v2-1, indicating a potential over-sensitivity to alignment mismatches. Such behavior highlights the limitations of these metrics in providing fair and unbiased evaluations across diverse model architectures.
VQA Score and TIFA Score are relative reliable metrics.
Among the evaluated metrics, VQA Score and TIFA Score stand out by assessing model performance on VQA tasks, rather than relying solely on subjective human preferences. This approach enhances the interpretability of the evaluation process. Additionally, we observed that the results from VQA Score and TIFA Score show a stronger correlation with other established benchmarks. Based on these advantages, we recommend prioritizing these two metrics for evaluation. However, it is important to note that their effectiveness is constrained by the limitations of the VQA models utilized in the evaluation.
A.3 Fairness analysis
We evaluate fairness by examining the model’s performance across different genders and races. Specifically, we calculate the average performance for each node and its associated child nodes within the taxonomy tree constructed for objects. For example, the node “females” includes child nodes such as “waitresses,” and their combined performance is considered in the analysis.
Gender.
Figure 11: Average performance scores of all models across different genders were evaluated using various metrics.
In gender, we observe a notable performance gap between females and males, as could be seen from Figure 11, Models are better at generating male concepts.
Race.
Figure 12: Average performance scores of all models across different races evaluated using various metrics.
There are also performance gaps in different races. From Figure 12, we found that ”white (person)” and ”black (person)” perform better than ”asian (person)”, ”Indian (amerindian)”, and ”Latin American”.
Appendix B Correlation of Generate Any Scene with other Text-to-Vision generation benchmarks
Figure 13: Correlation of Generate Any Scene with other popular Text-to-Vision generation benchmarks.
The Generate Any Scene benchmark uniquely relies entirely on synthetic captions to evaluate models. To assess the transferability of these synthetic captions, we analyzed the consistency in model rankings across different benchmarks [76, 34, 75]. Specifically, we identified the overlap of models evaluated by two benchmarks and computed the Spearman correlation coefficient between their rankings.
As shown in the figure 13, Generate Any Scene demonstrates a strong correlation with other benchmarks, such as Conceptmix [76] and GenAI Bench [34], indicating the robustness and reliability of Generate Any Scene’s synthetic caption-based evaluations. This suggests that the synthetic captions generated by Generate Any Scene can effectively reflect model performance trends, aligning closely with those observed in benchmarks using real-world captions or alternative evaluation methods.
Appendix C Details of Taxonomy of Visual Concepts
To construct a scene graph, we utilize three primary types of metadata: objects, attributes, and relations, which represent the structure of a visual scene. Additionally, scene attributes—which include factors like image style, perspective, and video time span—capture broader aspects of the visual content. Together, the scene graph and scene attributes form a comprehensive representation of the scene.
Our metadata is further organized using a well-defined taxonomy, enhancing the ability to generate controllable prompts. This hierarchical taxonomy not only facilitates the creation of diverse scene graphs, but also enables fine-grained and systematic model evaluation.
Objects.
To enhance the comprehensiveness and taxonomy of object data, we leverage noun synsets and the structure of WordNet [45]. In WordNet, a physical object is defined as ”a tangible and visible entity; an entity that can cast a shadow.” Following this definition, we designate the physical object as the root node, constructing a hierarchical tree with all 28,787 hyponyms under this category as the set of objects in our model.
Following WordNet’s hypernym-hyponym relationships, we establish a tree structure, linking each object to its primary parent node based on its first-listed hypernym. For objects with multiple hypernyms, we retain only the primary parent to simplify the hierarchy. Furthermore, to reduce ambiguity, if multiple senses of a term share the same parent, we exclude that term itself and reassign its children to the original parent node. This approach yields a well-defined and disambiguated taxonomy.
Attributes.
The attributes of a scene graph represent properties or characteristics associated with each object. We classify these attributes into nine primary categories. For color, we aggregate 677 unique entries sourced from Wikipedia [72]. The material category comprises 76 types, referenced from several public datasets [64, 4, 79]. The texture category includes 42 kinds from the Describable Textures Dataset [14], while the architectural style encompasses 25 distinct styles [78]. Additionally, we collect 85 states, 41 shapes, and 24 sizes. For human descriptors, we compile 59 terms across subcategories, including body type and height. Finally, we collect 465 common adjectives covering general characteristics of objects to enhance the descriptive richness of our scene graphs.
Relationships.
We leverage the Robin dataset [49] as the foundation for relationship metadata, encompassing six key categories: spatial, functional, interactional, social, emotional, and symbolic. With 10,492 relationships, the dataset provides a comprehensive and systematic repository that supports modeling diverse and complex object interactions. Its extensive coverage captures both tangible and abstract connections, forming a robust framework for accurate scene graph representation.
Scene Attributes.
In Text-to-Vision generation tasks, people mainly focus on creating realistic images and art from a text description [58, 55, 5]. For artistic styles, we define scene attributes using 76 renowned artists, 41 genres, and 126 painting styles from WikiArt [59], along with 29 common painting techniques. For realistic imagery, we construct camera settings attributes across 6 categories: camera models, focal lengths, perspectives, apertures, depths of field, and shot scales. The camera models are sourced from the 1000 Cameras Dataset [15], while the remaining categories are constructed based on photography knowledge and common prompts in Text-to-Vision generation tasks [7, 70]. To control scene settings, we categorize location, weather and lighting attributes, using 430 diverse locations from Places365 [41], alongside 76 weathers and 57 lighting conditions. For video generation, we introduce attributes that describe dynamic elements. These include 12 types of camera rig, 30 distinct camera movements, 15 video editing styles, and 27 temporal spans. The comprehensive scene attributes that we construct allow for the detailed and programmatic Text-to-Vision generation generation.
Appendix D Details of Overall Performance (Section 3)
D.1 Detailed experiment settings
  • $\bullet$
    For Text-to-Image generation, we select a range of open-source models, including those utilizing UNet backbones, such as DeepFloyd IF [1], Stable Diffusion v2-1 [56], SDXL [52], Playground v2.5 [35], and Wuerstchen v2 [51], as well as models with DiT backbones, including Stable Diffusion 3 Medium [16], PixArt-$\alpha$ [11], PixArt-$\Sigma$ [12], FLUX.1-schnell [33], FLUX.1-dev [33], and FLUX 1. Closed-source models, such as DaLL-E 3 [5] and FLUX1.1 PRO [33], are also assessed to ensure a comprehensive comparison. All models are evaluated at a resolution of 1024 × 1024 pixels.
  • $\bullet$
    For Text-to-Video generation, we select nine open-source models: ModelScope [67], ZeroScope [61], Text2Video-Zero [30], CogVideoX-2B [80], VideoCrafter2 [10], AnimateLCM [65], AnimateDiff [19], FreeInit [74], and Open-Sora 1.2 [84]. We standardize the frame length to 16 across all video models for fair comparisons.
  • $\bullet$
    For Text-to-3D generation, we evaluate five recently proposed models: SJC [66], DreamFusion [53], Magic3D [38], Latent-NeRF [44], and ProlificDreamer [69]. We employ the implementation and configurations provided by ThreeStudio [18] and generate videos by rendering from 120 viewpoints. To accelerate inference, we omit the refinement stage. For Magic3D and DreamFusion, we respectively use DeepFloyd IF and Stable Diffusion v2-1 as their 2D backbones.
Metrics.
Across all Text-to-Image generation, Text-to-Video generation, and Text-to-3D generation, we employ five widely used Text-to-Vision generation metrics to comprehensively assess model performance:
  • $\bullet$
    Clip Score: Assesses semantic similarity between images and text.
  • $\bullet$
    VQA Score and TIFA Score: Evaluate faithfulness by generating question-answer pairs and measuring answer accuracy from images.
  • $\bullet$
    Pick Score and ImageReward Score: Capture human preference tendencies.
We also use metrics in VBench [24] to evaluate Text-to-Video generation models on fine-grained dimensions, such as consistency and dynamics, providing detailed insights into video performance.
For Text-to-Video generation and Text-to-3D generation tasks:
  • $\bullet$
    We calculate Clip Score, Pick Score, and ImageReward Score on each frame, then average these scores across all frames to obtain an overall video score.
  • $\bullet$
    For VQA Score and TIFA Score, we handle Text-to-Video generation and Text-to-3D generation tasks differently:
    • $\circ$
      In Text-to-Video generation tasks, we uniformly sample four frames from the 16-frame sequence and arrange them in a 2 × 2 grid image.
    • $\circ$
      For Text-to-3D generation tasks, we render images at 45-degree intervals from nine different viewpoints and arrange them in a 3 × 3 grid.
This sampling approach optimizes inference speed without affecting score accuracy [39].
D.2 Detailed overall results
We evaluate Text-to-Image generation, Text-to-Video generation, and Text-to-3D generation models on Generate Any Scene. The detailed results of each model and each metric are shown in  Tabs. 5, 6, 7 and 8
Model clip score pick score vqa score tifa score image reward score
Playground v2.5 [35] 0.2581 0.2132 0.5734 0.2569 0.2919
Stable Diffusion v2-1 [56] 0.2453 0.1988 0.5282 0.2310 -0.9760
SDXL [52] 0.2614 0.2046 0.5328 0.2361 -0.3463
Wuerstchen v2 [51] 0.2448 0.2022 0.5352 0.2239 -0.3339
DeepFloyd IF XL [1] 0.2396 0.1935 0.5397 0.2171 -0.8687
Stable Diffusion 3 Medium [16] 0.2527 0.2027 0.5579 0.2693 -0.0557
PixArt-$\alpha$ [11] 0.2363 0.2050 0.6049 0.2593 0.1149
PixArt-$\Sigma$ [12] 0.2390 0.2068 0.6109 0.2683 0.0425
FLUX.1-dev [33] 0.2341 0.2060 0.5561 0.2295 0.1588
FLUX.1-schnell [33] 0.2542 0.2047 0.6132 0.2833 0.1251
FLUX1.1 PRO [33] 0.2315 0.2065 0.5744 0.2454 -0.0361
Dalle-3 [5] 0.2518 0.2006 0.6871 0.4249 0.3464
Table 5: Overall performance of Text-to-Image generation models over 10K Generate Any Scene prompts.
Model clip score pick score image reward score VQA score TiFA score
VideoCraft2 [10] 0.2398 0.1976 -0.4202 0.5018 0.2466
AnimateDiff [19] 0.2610 0.1959 -0.7301 0.5255 0.2208
Open-Sora 1.2 [84] 0.2259 0.1928 -0.6277 0.5519 0.2414
FreeInit [74] 0.2579 0.1950 -0.9335 0.5123 0.2047
ModelScope [67] 0.2041 0.1886 -1.9172 0.3840 0.1219
Text2Video-Zero [30] 0.2539 0.1933 -1.2050 0.4753 0.1952
AnimateLCM [65] 0.2450 0.1987 -0.5754 0.4816 0.2176
CogVideoX-2B [80] 0.2038 0.1901 -1.2301 0.4585 0.1997
ZeroScope [61] 0.2289 0.1933 -1.1599 0.4892 0.2388
Table 6: Overall performance of Text-to-Video generation models over 10k Generate Any Scene prompts.
Model subject consistency background consistency motion smoothness dynamic degree aesthetic quality imaging quality
VideoCraft2 0.9768 0.9688 0.9833 0.3556 0.5515 0.6974
AnimateDiff 0.9823 0.9733 0.9859 0.1406 0.5427 0.5830
Open-Sora 1.2 0.9964 0.9907 0.9973 0.0044 0.5235 0.6648
FreeInit 0.9581 0.9571 0.9752 0.4440 0.5200 0.5456
ModelScope 0.9795 0.9831 0.9803 0.1281 0.3993 0.6494
Text2Video-Zero 0.8471 0.9030 0.8301 0.9999 0.4889 0.7018
AnimateLCM 0.9883 0.9802 0.9887 0.0612 0.6323 0.6977
CogVideoX-2B 0.9583 0.9602 0.9823 0.4980 0.4607 0.6098
ZeroScope 0.9814 0.9811 0.9919 0.1670 0.4582 0.6782
Table 7: Overall performance of Text-to-Image generation models over 10k Generate Any Scene prompts with VBench metrics.
Model clip score pick score vqa score tifa score image reward score
ProlificDreamer [69] 0.2125 0.1940 0.5411 0.2704 -1.2774
Latent-NeRF [44] 0.2115 0.1910 0.4767 0.2216 -1.5311
DreamFusion-sd [53] 0.1961 0.1906 0.4421 0.1657 -1.5582
Magic3D-sd [38] 0.1947 0.1903 0.4193 0.1537 -1.6327
SJC [66] 0.2191 0.1915 0.5015 0.2563 -1.4370
DreamFusion-IF [53] 0.1828 0.1857 0.3872 0.1416 -1.9353
Magic3D-IF [38] 0.1919 0.1866 0.4039 0.1537 -1.8465
Table 8: Overall performance of Text-to-3D generation models over 10k Generate Any Scene prompts.
D.3 Case study: Pairwise fine-grained model comparison
Evaluating models using a single numerical average score can be limiting, as different training data often lead models to excel in generating different types of concepts. By leveraging the taxonomy we developed for Generate Any Scene, we can systematically organize these concepts and evaluate each model’s performance on specific concepts over the taxonomy. This approach enables a more detailed comparison of how well models perform on individual concepts rather than relying solely on an overall average score. Our analysis revealed that, while the models may achieve similar average performance, their strengths and weaknesses vary significantly across different concepts. Here we present a pairwise comparison of models across different metrics.
Figure 14: Stable Diffusion v2-1 vs. Stable Diffusion 3 Medium on average VQA Score in fine-grained categories.
Figure 15: PixArt-$\Sigma$ vs. Stable Diffusion 3 Medium on average VQA Score in fine-grained categories.
Figure 16: FLUX.1-schnell vs. Stable Diffusion 3 Medium on average VQA Score in fine-grained categories.
Figure 17: PixArt-$\Sigma$ vs. FLUX.1-schnell on average VQA Score in fine-grained categories.
Appendix E Details of Application 1: Self-Improving Models (Section 4)
E.1 Experiment details
E.1.1 Captions Preparation
To evaluate the effectiveness of our iterative self-improving Text-to-Vision generation model, we generated three distinct sets of 10k captions using Generate Any Scene, covering a sample complexity range from 3 to 12. These captions were programmatically created to reflect a spectrum of structured scene graph compositions, designed to challenge and enrich the model’s learning capabilities.
For comparative analysis, we leveraged the Conceptual Captions (CC3M) [9] dataset, a large-scale benchmark containing approximately 3.3 million image-caption pairs sourced from web alt-text descriptions. CC3M is renowned for its diverse visual content and natural language expressions, encompassing a wide range of styles, contexts, and semantic nuances.
To ensure fair comparison, we randomly sampled three subsets of 10k captions from the CC3M dataset, matching the Generate Any Scene-generated caption sets in size. This approach standardizes data volume while enabling direct performance evaluation. The diversity and semantic richness of the CC3M captions serve as a robust benchmark to assess whether Generate Any Scene-generated captions can match or exceed the descriptive quality of real-world data across varied visual contexts.
E.1.2 Dataset Construction and Selection Strategies
For the captions generated by Generate Any Scene, we employed a top-scoring selection strategy to construct the fine-tuning training dataset, using a random selection strategy as a baseline for comparison. Specifically, for each prompt, the model generated eight images. Under the top-scoring strategy, we evaluated the generated images using the VQA score and selected the highest-scoring image as the best representation of the prompt. This process yielded 10k top-ranked images per iteration, from which the top 25% (approximately 2.5k images) with the highest VQA scores were selected to form the fine-tuning dataset.
In the random selection strategy, one image was randomly chosen from the eight generated per prompt, and 25% of these 10k randomly selected images were sampled to create the fine-tuning dataset, maintaining parity in data size.
For the CC3M dataset, each prompt was uniquely paired with a real image. From the 10k real image-caption pairs sampled from CC3M, the top 25% with the highest VQA scores were selected as the fine-tuning training dataset. This ensured consistency in data size and selection criteria across all methods, facilitating a rigorous and equitable comparison of fine-tuning strategies.
E.1.3 Fine-tuning details
We fine-tuned the Stable Diffusion v1-5 using the LoRA technique. The training was conducted with a resolution of 512 $\times$ 512 for input images and a batch size of 8. Gradients were accumulated over two steps. The optimization process utilized the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, an $\epsilon$ value of $1\times 10^{-8}$, and a weight decay of $10^{-2}$. The learning rate was set to $1\times 10^{-4}$ and followed a cosine scheduler for smooth decay during training. To ensure stability, a gradient clipping threshold of 1.0 was applied. The fine-tuning process was executed for one epoch, with a maximum of 2500 training steps. For the LoRA-specific configurations, we set the rank of the low-rank adaptation layers and the scaling factor $\alpha$ to be 128.
After completing fine-tuning for each epoch, we set the LoRA weight to 0.75 and integrate it into Stable Diffusion v1-5 to guide image generation and selection for the next subset. For the CC3M dataset, images from the subsequent subset are directly selected.
In the following epoch, the fine-tuned LoRA parameters from the previous epoch are loaded and used to resume training on the current subset, ensuring continuity and leveraging the incremental improvements from prior iterations.
E.2 More results of fine-tuning models
Aside from our own test set and GenAI benchmark, we also evaluated our fine-tuned Text-to-Image generation models on the Tifa Bench (Figure 18), where we observed the same trend: models fine-tuned with our prompts consistently outperformed the original Stable Diffusion v1-5 and CC3M fine-tuned models.
Figure 18: Results for Application 1: Self-Improving Models. Average TIFA score of Stable Diffusion v1-5 fine-tuned with different data over TIFA Bench.
Appendix F Details of Application 2: Distilling limitations (Section 5)
F.1 Collecting hard concepts
We selected 81 challenging object concepts where Stable Diffusion v1-5 and DaLL-E 3 exhibit the largest gap in VQA Score. To determine the score for each concept, we calculated the average VQA score of the captions containing that specific concept. The full list of hard concepts is shown below:
  1. 1.
    cloverleaf
  2. 2.
    aerie (habitation)
  3. 3.
    admixture
  4. 4.
    webbing (web)
  5. 5.
    platter
  6. 6.
    voussoir
  7. 7.
    hearthstone
  8. 8.
    puttee
  9. 9.
    biretta
  10. 10.
    yarmulke
  11. 11.
    surplice
  12. 12.
    overcoat
  13. 13.
    needlepoint
  14. 14.
    headshot
  15. 15.
    photomicrograph
  16. 16.
    lavaliere
  17. 17.
    crepe
  18. 18.
    tureen
  19. 19.
    bale
  20. 20.
    jetliner
  21. 21.
    square-rigger
  22. 22.
    supertanker
  23. 23.
    pocketcomb
  24. 24.
    filament (wire)
  25. 25.
    inverter
  26. 26.
    denture
  27. 27.
    lidar
  28. 28.
    volumeter
  29. 29.
    colonoscope
  30. 30.
    synchrocyclotron
  31. 31.
    miller (shaper)
  32. 32.
    alternator
  33. 33.
    dicer
  34. 34.
    trundle
  35. 35.
    paddle (blade)
  36. 36.
    harmonica
  37. 37.
    piccolo
  38. 38.
    handrest
  39. 39.
    rundle
  40. 40.
    blowtorch
  41. 41.
    volleyball
  42. 42.
    tile (man)
  43. 43.
    shuttlecock
  44. 44.
    jigsaw
  45. 45.
    roaster (pan)
  46. 46.
    maze
  47. 47.
    belt (ammunition)
  48. 48.
    gaddi
  49. 49.
    drawer (container)
  50. 50.
    tenter
  51. 51.
    pinnacle (steeple)
  52. 52.
    pegboard
  53. 53.
    afterdeck
  54. 54.
    scaffold
  55. 55.
    catheter
  56. 56.
    broomcorn
  57. 57.
    spearmint
  58. 58.
    okra (herb)
  59. 59.
    goatsfoot
  60. 60.
    peperomia
  61. 61.
    ammobium
  62. 62.
    gazania
  63. 63.
    echinocactus
  64. 64.
    birthwort
  65. 65.
    love-in-a-mist (passionflower)
  66. 66.
    ragwort
  67. 67.
    spicebush (allspice)
  68. 68.
    leadplant
  69. 69.
    barberry
  70. 70.
    hamelia
  71. 71.
    jimsonweed
  72. 72.
    undershrub
  73. 73.
    dogwood
  74. 74.
    butternut (walnut)
  75. 75.
    bayberry (tree)
  76. 76.
    lodestar
  77. 77.
    tapa (bark)
  78. 78.
    epicalyx
  79. 79.
    blackberry (berry)
  80. 80.
    stub
  81. 81.
    shag (tangle)
F.2 Experiment details
We conducted targeted fine-tuning experiments on Stable Diffusion v1-5 to evaluate Generate Any Scene’s effectiveness in distilling model compositionality and learning hard concepts. For each task, we selected a dataset of 778 Generate Any Scene captions paired with images generated by DaLL-E 3. For compositionality, we selected multi-object captions from the existing dataset of 10k Generate Any Scene captions and paired them with the corresponding images generated by DaLL-E 3. To address hard concept learning, we first used Stable Diffusion v1-5 to generate images based on the 10k Generate Any Scene captions and identified the hard concepts with the lowest VQA scores. These concepts were then used to create a subset of objects, which we recombined into our scene-graph based captions with complexity levels ranging from 3 to 9. Finally, we used DaLL-E 3 to generate corresponding images for these newly composed captions.
The fine-tuning configurations were consistent with those used in the self-improving setup (Appendix E.1.3). To accommodate the reduced dataset size, the maximum training steps were set to 1000.
As a baseline, we randomly selected 778 images from 10k Generate Any Scene-generated images, using captions produced by Generate Any Scene. This ensured a controlled comparison between the targeted and random fine-tuning strategies.
Appendix G Details of Application 3: Generated content detector (Section 3)
G.1 Experiment details
In this section, our goal is to validate that the more diverse captions generated by Generate Any Scene can complement existing datasets, which are predominantly composed of real-world images paired with captions. By doing so, we aim to train AI-generated content detectors to achieve greater robustness.
Dataset preparation
We conducted comparative experiments between captions generated by Generate Any Scene and entries from the $D^{3}$ dataset. From the $D^{3}$ dataset, we randomly sampled 10k entries, each including a caption, a link to a real image, and an image generated by SD v1.4. Due to some broken links, we successfully downloaded 8.5k real images and retained 10k SD v1.4-generated images. We also used SD v1.4 to generate images based on 10k Generate Any Scene captions.
We varied the training data sizes based on the sampled dataset. Specifically, we sampled N real images from the 10k $D^{3}$ real images. For synthetic data, we compared N samples exclusively from $D^{3}$ with a mixed set of N/2 samples from 10k Generate Any Scene images and N/2 sampled from $D^{3}$, ensuring a total of N synthetic samples. Combined, this resulted in 2N training images. We tested 2N across various sizes, ranging from 2k to 10k.
Detector architecture and training
We employed ViT-T [73] and ResNet-18 [20] as backbones for the detection models. Their pretrained parameters on ImageNet-21k were frozen, and the final classification head was replaced with a linear layer using a sigmoid activation function to predict the probability of an image being AI-generated. During training, We used Binary Cross-Entropy (BCE) as the loss function, and the AdamW optimizer was applied with a learning rate of $2e^{-3}$. Training was conducted with a batch size of 256 for up to 50 epochs, with early stopping triggered after six epochs of no improvement in validation performance.
Testing
To evaluate the performance of models trained with varying dataset sizes and synthetic data combinations, we tested them on both GenImage and Generate Any Scene datasets to assess their in-domain and out-of-domain performance under different settings.
For GenImage, we used validation data from four models: SD v1.4, SD v1.5, MidJourney, and VQDM. Each validation set contained 8k real images and 8k generated images. For Generate Any Scene, we sampled 10k real images from CC3M and paired them with 10k generated images from each of the following models: Stable Diffusion v2-1, PixArt-$\alpha$, Stable Diffusion 3 Medium, and Playground v2.5. This created distinct test sets for evaluating model performance across different synthetic data sources.
Detector Data Scale (2N)
SDv1.4
(In-domain, same model)
SDv2.1 Pixart-$\alpha$ SDv3-medium Playground v2.5
Average
(In-domain, cross model)
$D^{3}$ + Ours $D^{3}$ $D^{3}$ + Ours $D^{3}$ $D^{3}$ + Ours $D^{3}$ $D^{3}$ + Ours $D^{3}$ $D^{3}$ + Ours $D^{3}$ $D^{3}$ + Ours $D^{3}$
Resnet-18 2k 0.6561 0.6663 0.7682 0.6750 0.7379 0.606 0.7509 0.6724 0.7380 0.5939 0.7488 0.6368
4k 0.6751 0.6812 0.7624 0.6853 0.7328 0.6494 0.7576 0.7028 0.7208 0.6163 0.7434 0.6635
6k 0.6780 0.6995 0.7886 0.6870 0.7493 0.6586 0.7768 0.7285 0.7349 0.6335 0.7624 0.6769
8k 0.6828 0.6964 0.7710 0.6741 0.7454 0.6418 0.7785 0.7186 0.7215 0.6033 0.7541 0.6595
10k 0.6830 0.6957 0.7807 0.6897 0.7483 0.6682 0.7781 0.7326 0.7300 0.6229 0.7593 0.6784
ViT-T 2k 0.6759 0.6672 0.7550 0.6827 0.7585 0.6758 0.7473 0.6941 0.7327 0.6106 0.7484 0.6658
4k 0.6878 0.6871 0.7576 0.7000 0.7605 0.7071 0.7549 0.7217 0.7221 0.6144 0.7488 0.6858
6k 0.6898 0.6891 0.7663 0.6962 0.7666 0.7164 0.7629 0.7238 0.7303 0.6134 0.7565 0.6875
8k 0.6962 0.6974 0.7655 0.6894 0.7712 0.7253 0.7653 0.7253 0.7381 0.6344 0.7600 0.6936
10k 0.6986 0.6984 0.7828 0.6960 0.7777 0.7275 0.7786 0.7334 0.7330 0.6293 0.7680 0.6966
Table 9: F1-Score Comparison of ResNet-18 and ViT-T Detectors Trained with $D^{3}$ and $D^{3}$+ Generate Any Scene Across In-Domain Settings
G.2 Results
Table 10 and Table 9 evaluate the performance of ResNet-18 and ViT-T detection backbones trained on datasets of varying sizes and compositions across in-domain (same model and cross-model) and out-of-domain settings. While models trained with $D^{3}$ and Generate Any Scene occasionally underperform compared to those trained solely on $D^{3}$ in the in-domain same-model setting, they exhibit significant advantages in both in-domain cross-model and out-of-domain evaluations. These results demonstrate that incorporating our data (Generate Any Scene) into the training process enhances the detector’s robustness. By supplementing existing datasets with Generate Any Scene under the same training configurations and dataset sizes, detectors achieve stronger cross-model and cross-dataset capabilities, highlighting improved generalizability to diverse generative models and datasets.
Detector Data Scale (2N) SDv1.5 VQDM Midjourney
Average
(Out-of-domain)
$D^{3}$ + Ours $D^{3}$ $D^{3}$ + Ours $D^{3}$ $D^{3}$ + Ours $D^{3}$ $D^{3}$ + Ours $D^{3}$
Resnet-18 2k 0.6515 0.6591 0.5629 0.5285 0.5803 0.5647 0.5982 0.5841
4k 0.6709 0.6817 0.5693 0.5428 0.6016 0.5941 0.6139 0.6062
6k 0.6750 0.6963 0.5724 0.5327 0.6084 0.6072 0.6186 0.6121
8k 0.6792 0.6965 0.5716 0.5282 0.6097 0.5873 0.6202 0.6040
10k 0.6814 0.6955 0.5812 0.5454 0.6109 0.6040 0.6245 0.6150
ViT-T 2k 0.6755 0.6685 0.5443 0.4966 0.6207 0.6066 0.6135 0.5906
4k 0.6845 0.6865 0.5591 0.4971 0.6416 0.6149 0.6284 0.5995
6k 0.6900 0.6890 0.5580 0.4948 0.6455 0.6259 0.6313 0.6032
8k 0.6940 0.6969 0.5553 0.4962 0.6495 0.6387 0.6329 0.6106
10k 0.6961 0.6988 0.5499 0.4975 0.6447 0.6358 0.6302 0.6107
Table 10: F1-Score Comparison of ResNet-18 and ViT-T Detectors Trained with $D^{3}$ and $D^{3}$+ Generate Any Scene Across Out-of-Domain Settings